# Development of a time correction algorithm for a precise synchronization of a free-running Rubidium atomic clock with the GPS Time

### v1

@@ 4 Discussion
Figure 17: Standard deviation of the residuals distributions between the Rb and the GPS Time after the offline (blue) or online (orange) correction as a function of the correction time window. Quadratic fits of the Septentrio data are used for the offline correction whereas linear fits are used for the online correction. The performance on simulated data is also shown for the $10560$ s time window with triangle markers.
Figure 17 shows the standard deviation of the Rb vs GPS Time difference after correction as a function of the correction time window’s width. The performance of the offline and online corrections on experimental data (colored dots) are compared to the performance we had obtained on simulated data (colored triangles) with a correction time window of $2880$ s, $28800$ s and $240000$ s. Note that these simulated data were only taking into account phase white noise, frequency white noise and frequency random walk components. No additional uncertainties were added to take into account other types of noise (e.g: flicker noise) or experimental conditions (e.g: imperfect calibrations, imperfect PHM time signal). These differences can explain the slightly better achievable performance obtained on simulated data ($0.64$-$1.15$ ns at $28,800$ s) compared to experimental data ($0.81$-$1.67$ ns at $28,800$ s) and the fact that the residuals are minimal with a time window of $28,800$ s with simulated data and $10,560$ s with experimental data. For both corrections, very similar performance of synchronization with GPS Time are obtained for correction time windows below $30,000$ s so there is no need to have much shorter windows. This result is consistent with the fact that, as seen in Figure 2, the stability of the Rubidium signal becomes limited by the frequency random walk for averaging times around $10^{4}$ s. It is also for similar averaging time windows that the Rubidium clock stability becomes worse than that of GPS Time. It was thus expected to find that similar correction time windows or shorter ones would be needed to efficiently correct for the random walk. The offline correction seems to provide a slightly better synchronization to GPS Time (down to $\sim 1$ ns) but the precision achievable with the online correction is already more than satisfying (better than $5$ ns for correction time windows below $100,000$ s) for synchronization between several experimental sites. Indeed, the needed level of synchronization is usually of the order of $100$ ns for those applications.
Note that the algorithms presented here were optimized to correct time stamps, provided by a free-running Rubidium clock, that were already corrected to eliminate the impact of the Rubidium clock frequency linear drift. This initial correction was done here once and for all by fitting the first few days of the time difference with respect to the GPS Time. This was found to be enough for a data-taking of around $50$ days. However, for a longer data-taking, one could regularly update this preliminary correction by fitting again the last few days of Septentrio data and subtracting the fit result before moving to the fit with shorter time windows (typically $10,560$ s) dedicated to the random walk correction.
@@ 5 Conclusions
In this paper, we presented a simple way to use time comparisons to GPS Time to synchronize the time stamps, generated using a free-running Rubidium clock, close to UTC while preserving its short term stability and correcting the long term frequency random walk. This method has the advantage of using relatively cheap instruments and to be applicable online for a real-time synchronization as well as to be robust against GPS signal reception failures. The online method could be applied for the real-time synchronization between several experimental sites in long-baseline neutrino physics experiments.
This method consists in fitting the GPS Time vs Rb measured by a GNSS receiver with a piece-wise polynomial function of time and in subtracting the result to the generated time stamps. The method was first designed and validated with simulated signals before assessing its performance on real data. We evaluated the performance of this correction by quantifying the stability of the clock signal before and after the correction using the Overlapping Allan Standard Deviation. We showed that the optimal length of the time window for the fit of the GPS Time vs Rb seats around $10,000$ seconds, corresponding to $11$ data points from the receiver. This time window allowed to maintain the best possible short term stability while correcting efficiently the frequency random walk. After correction with this time window, the difference to GPS Time stays within a window of ±$\pm 3.5$ ns (±$\pm 5$ ns) for the offline (resp. online) correction during the whole period of $\sim 50$ days of measurement. This performance largely meets the usual requirements for long-baseline neutrino physics experiments, like Hyper-Kamiokande and DUNE. Note that we do not expect the performance of the correction to be heavily degraded by isolated missing or outlier measurements from the receiver. However, this correction requires a constant monitoring of the Rubidium time signal with a GNSS receiver (or other reference that can be linked to UTC). One should thus make sure that such a reference is available in the long term and that there is no possibility to loose it for long periods (e.g.: several hours).
Acknowledgments

### v2

@@ 4 Discussion
Figure 15 shows the standard deviation of the Rb vs GPS Time difference after correction as a function of the correction time window’s width.
Figure 15: Standard deviation of the residuals distributions between the Rb and GPS Time after the offline (blue) or online (orange) correction as a function of the correction time window. Quadratic fits of the Septentrio data are used for the offline correction whereas linear fits are used for the online correction. The performance on simulated data is also shown for three values of the correction time windows.
The performance of the offline and online corrections on experimental data (colored dots) are compared to the performance we had obtained on simulated data (colored triangles) with a correction time window of $2880$ s, $28800$ s and $240000$ s. Note that these simulated data were only taking into account phase white noise, frequency white noise and frequency random walk components. In particular, the measured data also contain a linear frequency drift and this main difference could partly explain the difference of performances observed between data and simulation. Also, no additional uncertainties were added to take into account other types of noise (e.g: flicker noise) or experimental conditions (e.g: imperfect calibrations, etc.). For both corrections, very similar performance of synchronization with GPS Time are obtained for correction time windows below $30,000$ s so there is no need to have much shorter windows. This result is consistent with the fact that, as seen in Figure 6, the stability of the Rubidium signal becomes worse than that of the GPS around $10^{4}$ s. The offline correction seems to provide a slightly better synchronization to GPS Time (down to $\sim 0.3$ ns) but the precision achievable with the online correction is already more than satisfying: better than $5$ ns for correction time windows below $100,000$ s.
@@ 5 Conclusions
In this paper, we presented a simple way to use time comparisons to GPS Time to synchronize the time stamps, generated using a free-running Rubidium clock, close to UTC while preserving its short term stability and correcting for the long term frequency random walk and deterministic drift. This method has the advantage of using relatively cheap instruments and to be applicable online for a real-time synchronization as well as to be robust against punctual GPS signal reception failures. The online method could be applied for the real-time synchronization between several experimental sites in long-baseline accelerator neutrino experiments as well as for other detectors involved in multi-messenger astrophysics measurements.
The proposed method consists in fitting GPS Time vs Rubidium measured by a GNSS receiver with a piece-wise polynomial function of time and in subtracting the result to the generated time stamps. The method was first designed and validated with simulated signals before assessing its performance on real data. We evaluated the performance of this correction by quantifying the stability of the clock signal before and after the correction using the Overlapping Allan Standard Deviation. We showed that the optimal length of the time window for the fit of GPS Time vs Rubidium seats around $10,000$ seconds, corresponding to around $10$ data points from the receiver. This time window allowed to maintain the best possible short term stability while correcting efficiently the frequency random walk. After correction with this time window, the difference to GPS Time stays within a window of ±$\pm 5$ ns for both offline and online corrections during the whole period of $35$ days of measurement. This performance largely meets the usual requirements for long-baseline accelerator neutrino experiments, like Hyper-Kamiokande and DUNE. Note that we do not expect the performance of the correction to be heavily degraded by isolated missing or outlier measurements from the receiver. However, this correction requires a constant monitoring of the Rubidium time signal with a GNSS receiver (or other reference that can be linked to UTC). One should thus make sure that such a reference is available in the long term and that there is no risk of loosing it for long periods (e.g.: several hours).
Fundings: